Dinklage Corp. has 8 million shares of common stock outstanding. The current share price is $74, and the book value per share is $5. The company also has two bond issues outstanding. The first bond issue has a face value of $80 million, has a 9 percent coupon, and sells for 95 percent of par. The second issue has a face value of $60 million, has a 10 percent coupon, and sells for 108 percent of par. The first issue matures in 24 years, the second in 8 years.

  

The most recent dividend was $4.6 and the dividend growth rate is 5 percent. Assume that the overall cost of debt is the weighted average of that implied by the two outstanding debt issues. Both bonds make semiannual payments. The tax rate is 35 percent.

  

What is the company's WACC?

Multiple Choice

  • 10.45%

  • 9.53%

  • 8.6%

  • 18.13%

  • 9.07%
Answer #1
As per DDM
Price = recent dividend* (1 + growth rate )/(cost of equity - growth rate)
74 = 4.6 * (1+0.05) / (Cost of equity - 0.05)
Cost of equity% = 11.53
MV of equity=Price of equity*number of shares outstanding
MV of equity=74*8000000
=592000000
MV of Bond1=Par value*bonds outstanding*%age of par
MV of Bond1=1000*80000*0.95
=76000000
MV of Bond2=Par value*bonds outstanding*%age of par
MV of Bond2=1000*60000*1.08
=64800000
MV of firm = MV of Equity + MV of Bond1+ MV of Bond 2
=592000000+76000000+64800000
=732800000
Cost of debt
Bond1
                  K = Nx2
Bond Price =∑ [(Semi Annual Coupon)/(1 + YTM/2)^k]     +   Par value/(1 + YTM/2)^Nx2
                   k=1
                  K =24x2
950 =∑ [(9*1000/200)/(1 + YTM/200)^k]     +   1000/(1 + YTM/200)^24x2
                   k=1
YTM1 = 9.533
Bond2
                  K = Nx2
Bond Price =∑ [(Semi Annual Coupon)/(1 + YTM/2)^k]     +   Par value/(1 + YTM/2)^Nx2
                   k=1
                  K =8x2
1080 =∑ [(10*1000/200)/(1 + YTM/200)^k]     +   1000/(1 + YTM/200)^8x2
                   k=1
YTM2 = 8.6
Firm cost of debt=YTM1*(MV bond1)/(MV bond1+MV bond2)+YTM2*(MV bond2)/(MV bond1+MV bond2)
Firm cost of debt=9.533*(76000000)/(76000000+64800000)+8.6*(76000000)/(76000000+64800000)
Firm cost of debt=9.1%
After tax cost of debt = cost of debt*(1-tax rate)
After tax cost of debt = 9.1*(1-0.35)
= 5.915
Weight of equity = MV of Equity/MV of firm
Weight of equity = 592000000/732800000
W(E)=0.8079
Weight of debt = MV of Bond/MV of firm
Weight of debt = 140800000/732800000
W(D)=0.1921
WACC=after tax cost of debt*W(D)+cost of equity*W(E)
WACC=5.92*0.1921+11.53*0.8079
WACC% = 10.45